### v1

# Audio Visual Segmentation Through Text Embeddings

Abstract
The goal of Audio-Visual Segmentation (AVS) is to localize and segment the sounding source objects from the video frames. Researchers working on AVS suffer from limited datasets because hand-crafted annotation is expensive. Recent works attempt to overcome the challenge of limited data by leveraging the segmentation foundation model, SAM, prompting it with audio to enhance its ability to segment sounding source objects. While this approach alleviates the model’s burden on understanding visual modality by utilizing pre-trained knowledge of SAM, it does not address the fundamental challenge of the limited dataset for learning audio-visual relationships. To address these limitations, we propose AV2T-SAM, a novel framework that bridges audio features with the text embedding space of pre-trained text-prompted SAM. Our method leverages multimodal correspondence learned from rich text-image paired datasets to enhance audio-visual alignment. Furthermore, we introduce a novel feature, , which emphasizes shared semantics of audio and visual modalities while filtering irrelevant noise. Experiments on the AVSBench dataset demonstrate state-of-the-art performance on both datasets of AVSBench. Our approach outperforms existing methods by effectively utilizing pretrained segmentation models and cross-modal semantic alignment.
Index Terms— Audio-Visual Segmentation, SAM
1 Introduction
Audio-visual segmentation (AVS) is the task of segmenting sounding source objects from the video frames. Without pixel-level supervision, early researchers tackled the sounding source localization (SSL) task by leveraging self-supervised learning [1, 2] , exploiting the semantic alignment between image and audio pair data. However, this approach results in coarse localization of sounding objects, which limits the application in fields that require fine-grained segmentation masks. Zhou et al. [3] introduced the AVSBench dataset, annotated with a pixel-level segmentation map. The initial works on this dataset often fuse audio and image embeddings [3, 4] and then decode the segmentation mask from the fused audio-visual embeddings.
The Segment Anything Model (SAM) [5, 6] introduces an innovative approach to segmentation tasks by leveraging large-scale pretraining, similar to the way large language models use prompt-based learning [7]. The prompting idea has been applied to AVS. Mo and Tian [8] pioneered the idea of audio-prompted SAM by providing audio as a prompt to the SAM decoder. Liu et al. [9] advanced the audio-prompted SAM by introducing adapters into the image encoder of SAM. Audio-prompted SAM achieved better performance on the AVS task by leveraging the image segmentation foundation model, which mitigates the problem of the limited dataset. However, while SAM empowers the capacity of image understanding, the burden of learning audio-visual correspondence persists due to the scarcity of labeled data.
To address the data scarcity issue, we propose AV2T-SAM, which projects audio prompts into a text embedding space. This enables the use of knowledge from pre-trained text-prompted SAM models. Text-image models benefit from a wealth of text-image paired datasets, surpassing the availability of audio-visual data. By transforming audio inputs into a format compatible with pre-trained text-prompted SAM, AV2T-SAM not only capitalizes on SAM’s segmentation expertise but also enhances the learning of audio-visual correspondence through access to the text embedding space.
Specifically, we introduce a novel feature, , which aligns audio and visual embeddings into a shared semantic space, leveraging the pre-trained cross-modal encoders CLIP [10] and CLAP [11]. By leveraging modality-specific projections, the model combines these embeddings through element-wise multiplication, capturing the intersection of audio and visual modalities. This design emphasizes shared semantics while reducing irrelevant information, producing a more robust and contextually enriched representation. Through this approach, our model effectively enhances audio-visual correspondence, enabling more accurate and reliable segmentation performance. Also, we report a vision bias problem in the single sound source dataset (S4). We found the vision bias from the dataset by performing better than previous state-of-the-art models without using audio information.
In summary, our contributions are as follows:
We propose AV2T-SAM (Audio-Visual to Text SAM), the novel framework of utilizing a pre-trained text-prompted SAM by projecting audio features into the text embedding space. This strategy enables us to effectively harness the rich semantic information embedded in large-scale pre-trained cross-modal encoders, addressing the limitations posed by the scarcity of audio-visual segmentation masks.
We introduce a feature which captures the intersection of audio and visual modalities, emphasizing shared semantics and filtering out irrelevant information, resulting in a more contextually enriched and discriminative feature for segmentation tasks.
We demonstrate that the single sound source (S4) dataset contains a serious vision bias problem by surpassing previous state-of-the-art approaches without using any audio information.
2 Related works
2.1 Multimodal Learning with Segment Anything Model
Segment Anything Model (SAM) [5] is the first foundation model for image segmentation tasks with various prompts. The model is pretrained with more than 11M images and 1B masks. Many studies verified capability of SAM on diverse downstream tasks, including medical image segmentation [12] and pose estimation [13].
Beyond vision tasks, SAM is also widely used in multimodal learning due to its flexible prompt interactions. Wang et al.[14] employ remote sensing data to prompt SAM, while Liu et al.[9] use audio to segment sounding objects. Researchers have also explored text-based prompting for SAM [15]. In this study, we introduce applying a pretrained text-prompted SAM on audio-visual segmentation.
2.2 Audio-Visual Segmentation
With the trends of multimodal learning, audio-visual tasks also got much attention. Aiming for better audio-visual correspondence, researchers worked on Sound Source Localization task [16, 17] that aims to localize the sound source portion from the frame. Due to the lack of a labeled dataset, researchers had to rely on Self-Supervised learning, which, while effective in leveraging unlabelled data, comes with limited applications to fine-grained tasks such as segmentation. Zhou et al. [3] introduce Audio-Visual Segmentation with a labeled AVSBench dataset. The AVSBench dataset enabled researchers to work on the Audio-Visual Segmentation task, which is more fine-grained than the localization task. With the new dataset, researchers started exploring audio-visual segmentation [3, 18]. Gao et al. [4] proposed an architecture that leverages the transformer in the AVS task. Liu et al. [19] introduced a new framework that separates the segmentation process and the verifying sounding object process.
Mo and Tian [8] first explored utilizing the segment anything model with an audio prompt. Claiming the two-layer SAM decoder, which was previously prompted, is too shallow to model the audio-visual correspondence, Liu et al. [9] introduced the adapter that introjects audio information to the frozen SAM encoder. Seon et al. [20] employed temporal information which was not considered on sam-based models. Nguyen and Park [21] advanced the adapter mechanism by introducing an additional trainable layer to each block of the SAM encoder and utilized the residual audio encoder. In this work, we will exploit the text embedding space and the power of pretrained text-prompted SAM [15] to advance the performance of SAM [5] on the AVS.
3 Method
3.1 Problem Formation
The input video of audio-visual segmentation (AVS) comprises of a series of visual frames and corresponding audio. The image frames are represented as V = , where () where denotes the -th visual frame with spatial dimensions , and is the total number of frames in the video. Each frame is paired with a corresponding 1-second audio waveform , where represents the number of audio samples in 1 second. The goal of the AVS task is to segment all sound source objects from each frame. The ground truth segmentation map is provided by binary masks where
3.2 Text-prompted Segmentation Module
The pre-trained text-prompted SAM model is used in our framework as a backbone to leverage its robust text-visual correspondence knowledge, which is from extensive text-image datasets [22, 23, 24]. While this module is interchangeable, we particularly used EVF-SAM [15]. The model employs the multimodal encoder, which is more efficient than adopting the Large Language model and performs better than the simple text encoder as the CLIP. While the original EVF-SAM model uses text as input, our method does not rely on explicit text. Instead, we map audio-visual features into the language space. To achieve this, we adapt their methods by removing the text encoder but keeping the multimodal encoder fixed, leveraging its strong adaptability to prompt SAM with text embeddings effectively.
3.3 Projection Module
This module consists of 2 Multi-layer Perceptron (MLP) projection layers that connect feature to the text embedding space required to prompt the pre-trained text-prompted segmentation model.
3.4 Semantically Aligned Features
To enhance the semantic alignment between audio and visual modalities, we propose feature. Let and denote the normalized embeddings for audio and visual inputs, respectively, obtained from CLIP[10] and CLAP[11] encoders. These embeddings are projected into a shared embedding space of dimension using learned projection matrices and . The projected embeddings are computed as follows:
where and represent the transformed embeddings in the shared space. To capture the intersection of semantic information from both modalities, we calculate:
, where denotes the Hadamard (element-wise) product, and is the fused feature representation.
This operation ensures that the resulting feature emphasizes shared semantics between the two modalities while filtering out modality-specific noise. The design leverages the complementary nature of audio and visual information, providing a robust and contextually enriched representation that enhances downstream segmentation tasks. By projecting both modalities into a unified semantic space and fusing them through intersection, our approach effectively aligns and refines cross-modal information, addressing challenges in audio-visual correspondence learning.
3.5 Adapters
The mask decoder of SAM, which is two transformer layers, is too shallow to fuse the audio and visual information [9]. However, fine-tuning the image encoder of SAM requires considerable computing resources. To efficiently fuse audio and visual information while not entirely fine-tuning the image encoder, Liu et al. [9] suggested an audio adapter mechanism that trains only the adapter module which injects audio information to the frozen image encoder.
In this work, we employ the adapter module in the mask decoder of SAM with to effectively fuse prompt information to the image encoder. Specifically, for the -th layer of a transformer block, the -th adapter projects the prompt information to the output of the transformer block and repeats the prompt information along the spatial dimensions to match the dimensions of the image features, denoted as . The adapted features are then added to the output of the previous encoder layer, , to form the input to the -th layer, , as follows:
This approach ensures the efficient fusion of audio and visual information while preserving computational efficiency by freezing the image encoder during training.
3.6 Learning Objectives
For training the model, we combined the Binary Cross-Entropy (BCE) loss and the Intersection over Union (IoU) loss, following the Liu et al. [9].
4 Experimental Setup
4.1 Dataset
AVSBench [3] is the pixel-level audio-visual segmentation dataset, which consists of 5-second videos downloaded from YouTube. AVSBench contains two datasets: Single Sound Source Dataset (S4) and Multi Sound Source Dataset (MS3).
Single Sound Source dataset (S4). This dataset includes videos with only a single sound source. The dataset contains 4,932 videos in total over 23 categories and is split into training, validation, and test sets with 3,452, 740, and 740 samples, respectively. Notably, annotation details vary across subsets: only the first frame of each training video is annotated with binary masks, while the validation and test sets include full annotations for all five frames of every video.
Multi Sound Sources dataset (MS3). The dataset contains 424 videos and is split into training, validation, and test sets with 296, 64, and 64 samples, respectively. For all splits, including training, every frame is annotated with the segmentation mask.
4.2 Evaluation Metrics
Following the original work [3], we use the mean Intersection over Union () and F-score () to evaluate our model. The computes the average IoU of predicted mask and ground truth masks over total frames, and computes the harmonic mean of precision and recall.
4.3 Implementation Details
We utilized EVF-SAM2 [15], which integrates SAM-2-L [6] as the SAM backbone and Beit-3-L [25] as the multimodal encoder for receiving a text prompt. For experiments using the SAM1 [5] as a backbone, we employed EVF-SAM1, which consists of SAM-H and Beit-3-L. The input image resolution was set to , and We trained the model for 40 epochs using the AdamW optimizer.
5 Experiments
5.1 Main Results
| Category | Methods | Adapter | S4 (V1S) | MS3 (V1M) | ||
| AVS | AVSBench | - | 78.74 | 0.879 | 54.00 | 0.645 |
| ECMVAE | - | 81.74 | 0.901 | 57.84 | 0.708 | |
| AVSegFormer | - | 82.06 | 0.899 | 58.36 | 0.693 | |
| AVSC | - | 81.29 | 0.886 | 59.50 | 0.657 | |
| AuTR | - | 80.40 | 0.891 | 56.20 | 0.672 | |
| GAVS | - | 80.06 | 0.902 | 63.70 | 0.774 | |
| AQFormer | - | 81.60 | 0.894 | 61.10 | 0.721 | |
| COMBO | - | 84.70 | 0.919 | 59.20 | 0.712 | |
| SAM | AV-SAM | ✗ | 40.47 | 0.566 | - | - |
| AP-SAM | ✗ | 69.61 | 0.796 | 51.58 | 0.578 | |
| SAMA-AVS | ✓ | 81.53 | 0.886 | 63.14 | 0.691 | |
| SAVE | ✓ | 85.11 | 0.912 | 67.01 | 0.739 | |
| ST-BAVA | ✓ | 82.46 | 0.906 | 69.01 | 0.776 | |
| AV2T-SAM | ✗ | 83.11 | 0.901 | 55.81 | 0.634 | |
| AV2T-SAM | ✓ | 85.78 | 0.919 | 65.76 | 0.738 | |
| SAM2 | AV2T-SAM | ✗ | 85.63 | 0.920 | 64.47 | 0.704 |
| AV2T-SAM | ✓ | 86.67 | 0.924 | 69.65 | 0.777 | |
We compare our model with state-of-the-art methods in Table 1. Our proposed model achieved new state-of-the-art performance on both subsets of ASVBench. On the S4 dataset, our model outperform the previous state-of-the-art, SAVE [21], with an improvement of 1.56 without introducing extra adapter layers into every image encoder block. Also, on the MS3 dataset, we exceed the previous state-of-the-art, ST-BAVA [20], without utilizing temporal information, which enabled ST-BAVA to outperform other approaches significantly.
For a more fair comparison, we also run experiments with the model with SAM1[26] backbone, considering all previous SAM-based approaches employ SAM1 as a backbone. Under this setting, our model continues to surpass all existing state-of-the-art methods on the S4 dataset. On the MS3 dataset, it falls marginally short of the best result but still maintains competitive performance. When using the SAM2 [6] backbone, many complex components, such as adapters within every block of the image encoder as in SAVE, and a module for temporal understanding as in ST-BAVA, are unnecessary.
Figure 2 compares our segmentation results with those of existing models. In the left portion of the figure, we show one example of the results on the S4 dataset. TPAVI [3] captures only part of the guitar, suggesting weak visual object understanding. SAMA-AVS [9] leverages SAM to accurately segment the guitar but struggles to separate it from nearby objects. In contrast, our model fully segments the guitar and distinguishes it from its surroundings, made possible by strong audio-visual correspondence guided by the pretrained text-prompted SAM. A similar pattern emerges in the right portion of the figure, which depicts results on the MS3 dataset. Again, TPAVI exhibits partial segmentation of objects. While SAMA-AVS identifies the correct objects, it fails to separate the keyboard from a person. Our model, however, successfully segments the keyboard and isolates it from other objects. These findings underscore our model’s superior ability to detect and correlate audio with its corresponding visual region, thanks to the capabilities of the pre-trained text-prompted SAM.
5.2 Ablation on Adapters
We present an ablation study of our framework in Table 1. In general, methods without adapters perform not as well as those with adapters, and this also applies to our method.
Interestingly, comparing all the methods that do not use adapters, our proposed method still achieves much better performance than the previous method without adapters. Moreover, compared to existing methods that rely on adapters, our approach delivers comparable results—an outcome not achieved by any prior methods. This highlights the effectiveness of our framework and feature design.
From the result, we can notice that our framework, even without the adapter module, enhances performance dramatically on S4, the dataset which requires shallow audio-visual fusion. However, the model without the adapter module still suffers on the Multi Sound Sources dataset (MS4), the dataset that requires more profound audio-visual. This observation shows that while our approach strengthens audio-visual correspondence, it cannot replace the adapter module.
5.3 Ablation on Feature
In this section, we compare the model’s performance with different prompt features. As denoted in a Table. 2, we can observe that perform best on both metrics over both datasets. This implies that feature, which captures information from both audio and visual modalities, contains richer information as the prompt than just the CLAP feature, which contains only audio information.
It is worth noticing that the model prompted with CLIP [10] features surpasses the previous state-of-the-art on the S4 dataset, achieving a score of 86.67 on . Given that CLIP lacks any audio-related information, this result suggests that a performance of 86.67 on can be attained using only visual cues. This finding highlights a significant vision bias in the dataset, raising concerns about its effectiveness in evaluating true audio-visual understanding.
6 Conclusion
In this work, we have proposed a novel framework that leverages text modality embeddings and pre-trained text-prompted SAM on the audio-visual segmentation task. Also, we introduce a new feature that is aware of the intersection of audio and visual modalities for rich semantic information. Comprehensive experiment results have demonstrated that our approach performs comparable on the AVS task to state-of-the-arts models. Finally, we demonstrate the vision bias issue of the S4 dataset by surpassing previous state-of-the-art performance without using audio information.